# Advancing Universal Deep Learning for Electronic-Structure Hamiltonian Prediction of Materials

### v2

@@ A.1 Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
A.2 Introduction of Reciprocal Space Electronic-Structure Hamiltonians into Deep Learning Paradigm
@@
The self-consistent Hamiltonian $\mathbb{H}^{(T)}$, obtained through the procedure described in Section A.1, is inherently defined in real space. Its matrix elements $H_{\alpha\beta}^{(T)}$ are constructed over localized atomic orbital basis functions centered at atoms, and are truncated beyond a spatial cutoff. While real-space representations are efficient for representing local interactions, many physical phenomena such as band structures, effective low-energy models, and quasiparticle dynamics are most naturally described in reciprocal space.
@@
where $\theta$ is the angle between the eigenspaces corresponding to $\varepsilon_{n\mathbf{k}}$ and $\hat{\varepsilon}_{n\mathbf{k}}$. The theorem(Golub & Van Loan, 2013) highlights that due to the non-orthogonality of the orbital basis set, the errors in band energies and wave functions can be amplified by the condition number factor $\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}$. As a result, even a small error may cause the band eigenvalues and wave functions to deviate significantly from the true results, manifesting as the appearance of ghost states in the band structure.
@@
A.3 Details on Training Loss Functions
@@
where $\mu$ is a scalar coefficient, $\mathbf{S}(\mathbf{R})$ denotes the real-space overlap matrix, and $\mathbf{H}^{(0)}(\mathbf{R})$ denotes the real-space zeroth-step Hamiltonian matrix. Following the gauge-error formulation of Wang et al. (2024b), adding a shift term $\mu\cdot\mathbf{S}(\mathbf{R})$ to the Hamiltonian leaves all down-stream physical observables unchanged. In practice, $\mu$ is chosen as the solution that minimizes the overall loss, as established in Wang et al. (2024b). This removes the gauge freedom inherent in the Hamiltonian representation, facilitating more stable and efficient convergence of the neural network.
@@
$\begin{split}\mathrm{loss}(\mathbf{k})&=\mathbb{E}_{\mathbf{k}}\Big[\lambda_{\mathcal{P}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PP}(\mathbf{k}),\widetilde{\mathbf{H}}_{PP}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{\mathcal{Q}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{QQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{QQ}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{PQ}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{PQ}^{gt}(\mathbf{k},\mu)\big)\Big].\end{split}$ (21)
where $\lambda_{\mathcal{P}},\lambda_{\mathcal{Q}},\lambda_{PQ}$ are tunable hyperparameters controlling the relative importance of the three terms.
@@
where the value of $\mu$ is determined by $\frac{\partial\mathrm{loss}_{all}}{\partial\mu}=0$. It can be solved analytically as:
@@ A.4 Dataset Details
The crystal structures were obtained from the Materials Project database, from which a total of approximately 17,000 nonmagnetic compounds were randomly selected. Among them, 12,000 structures were used for training, 2,000 for validation, and 3,000 for testing. The statistical distributions of atomic species and atomic counts in the training, validation, and test sets are illustrated in Figures 4 and 5. Furthermore, the occurrence frequencies of different elements across the three subsets are presented in Figures 6(a)6(c).
@@ A.5 Implementation Details
For the input of the neural network, we adopt a cutoff radius of $8.0\,\text{\AA }$ to define the neighboring range in the atomic graph. The angular relations between atoms are represented using spherical harmonics with degrees $0\leq l\leq 5$, while the interatomic distances are encoded through a Gaussian basis expansion (Gong et al., 2023) with a preset base number of 64. The Transformer network consists of four stacked basic blocks. Each block contains an E(3)-symmetry layer normalization module, an E(3)-symmetry feed-forward module, an E(3)-symmetry multi-head graph attention module, and a TraceGrad module. The internal node features $\mathbf{f}^{\text{(node)}}_{a}$ and edge features $\mathbf{f}^{\text{(edge)}}_{ab}$, $\mathbf{f}^{\prime\text{(edge)}}_{ab}$, as well as $\mathbf{o}^{\text{(edge)}}_{ab}$, are represented in direct-sum state form with the following structure:
$32\times 0e+16\times 1e+16\times 1o+8\times 2e+8\times 2o+8\times 3e+8\times 3o+8\times 4e+8\times 4o+1\times 5o+1\times 6e,$
where the coefficients preceding $\times$ denote multiplicities, and the subsequent terms indicate tensor degrees together with their corresponding even (e) or odd (o) parities. For the TraceGrad module, the constructed SO(3)-invariant feature $z^{\text{(edge)}}_{ab}$ has a dimension of $256$. On the output side, to map the network outputs from the direct-sum E(3)-symmetric tensors into Hamiltonian matrices, we employ the conversion modules provided by Gong et al. (2023), thereby ensuring the exact symmetry of the predicted results with SU(2) symmetry. We employ an ensemble of four sub-models to predict the electronic-structure Hamiltonian. The first sub-model is responsible for predicting the Hamiltonian submatrices formed by atomic pairs with interatomic distances in the range $0,1.0\,\text{Å})$, where the case of distance equal to zero corresponds to the on-site Hamiltonian (i.e., the Hamiltonian formed by an atom with itself). The second sub-model handles atomic pairs with distances in the range $1.0\,\text{Å},2.0\,\text{Å})$, the third sub-model covers the range $2.0\,\text{Å},4.0\,\text{Å})$, and the fourth sub-model addresses the range $4.0\,\text{Å},6.0\,\text{Å})$. For atomic pairs with distances greater than 6.0 Å, we found that their self-consistent Hamiltonian is almost identical to the zeroth-step Hamiltonian numerically. Therefore, for these distant atoms, we bypass the neural network correction step and use the zeroth-step Hamiltonian as the final result.
@@
A.6 Band Structure Results
@@
A.7 Efficiency Comparison between Our Method and DFT
@@ A.8 Ablation Studies
Ablation@TraceGrad: In this ablation term, we remove the TraceGrad mechanism. Concretely, the supervision from the trace quantity is omitted in the loss function, and the gradient-based mechanism that delivers non-linearity from SO(3)-invariant features $z^{(\text{edge})}_{ab}$ to induce SO(3)-equivariant features via $\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{\prime\text{(edge)}}_{ab}}$ is also discarded.
@@
We train all of the ablation terms under the same number of epochs, optimizer, and scheduler as the full method (see Appendix A.5), then evaluate them on the testing set. The $R$-space errors are summarized in Table 4. Beyond $R$-space, because $k$-space is directly tied to downstream quantities (e.g., band structures), we visualize band predictions for Ablation@Loss-k (R-space only), Ablation@Loss-PQ (setting $\lambda_{PQ}=0$) versus the Full Method in Fig. 14.
From Table 4, the Full Method achieves the lowest Gauge MAE. The Full Method reduces the error by $17.6\%$, $52.3\%$, $20.7\%$, $23.8\%$, $12.2\%$, and $5.2\%$ compared with Ablation@Input, Ablation@Output, Ablation@TraceGrad, Ablation@Ensemble, Ablation@Loss-k, and Ablation@Loss-PQ, respectively. As shown in Fig. 14, Ablation@Loss-k, which removes the $k$-space supervision and relies solely on real-space loss, produces band structures with frequent ghost states: in many cases, while most $k$-points are predicted reasonably well, some $k$-points exhibit abrupt and severe deviations from the ground truth—hallmarks of non-physical artifacts. This phenomenon mainly arises from the error amplification effect analyzed in Appendix A.2, where the large condition number of the overlap matrix can magnify small real-space errors into significant $k$-space deviations. Importantly, such sparse but catastrophic failures cannot be effectively captured by real-space loss alone. Ablation@Loss-PQ, which augments the training with $k$-space supervision on the intra-subspace blocks ($P$ and $Q$), demonstrates better performance than Ablation@Loss-k, but still fails to completely suppress ghost states. The reason is that unphysical couplings between the low-energy subspace $P$ and the high-energy subspace $Q$ remain unpenalized, and these couplings are precisely what give rise to unphysical artifacts in the band structures. In contrast, the Full Method introduces an important penalty on the $PQ$ cross block, which has clear physical significance: for the exact Hamiltonian, $P$ and $Q$ are strictly decoupled, and any spurious $PQ$ couplings in the predicted Hamiltonian are the direct source of ghost states. By explicitly enforcing this decoupling, the $PQ$ loss term addresses the root cause of the artifacts. As a result, the full method produces band structures in excellent agreement with first-principles DFT and free of ghost states. This comparison clearly demonstrates the necessity of our $k$-space loss design, in particular the $PQ$ penalty, for ensuring the physical reliability of predicted band structures.

### v3

@@ E.2 Initial Edge Features
Appendix F Introduction of Reciprocal Space Electronic-Structure Hamiltonians into Deep Learning Paradigm
@@
The self-consistent Hamiltonian $\mathbb{H}^{(T)}$, obtained through the procedure described in Section A, is inherently defined in real space. Its matrix elements $H_{\alpha\beta}^{(T)}$ are constructed over localized atomic orbital basis functions centered at atoms, and are truncated beyond a spatial cutoff. While real-space representations are efficient for representing local interactions, many physical phenomena such as band structures, effective low-energy models, and quasiparticle dynamics are most naturally described in reciprocal space.
@@
where $\theta$ is the angle between the eigenspaces corresponding to $\varepsilon_{n\mathbf{k}}$ and $\hat{\varepsilon}_{n\mathbf{k}}$. The theorem (Golub and Van Loan, 2013) highlights that due to the non-orthogonality of the orbital basis set, the errors in band energies and wave functions can be amplified by the condition number factor $\frac{\kappa(S(\mathbf{k}))}{\|S(\mathbf{k})\|_{2}}$. As a result, even a small error may cause the band eigenvalues and wave functions to deviate significantly from the true results, manifesting as the appearance of ghost states in the band structure.
@@
Appendix G Details on Training Loss Functions
@@
where $\mu$ is a scalar coefficient, $\mathbf{S}(\mathbf{R})$ denotes the real-space overlap matrix, and $\mathbf{H}^{(0)}(\mathbf{R})$ denotes the real-space zeroth-step Hamiltonian matrix. Following the gauge-error formulation of Wang et al. (2024c), adding a shift term $\mu\cdot\mathbf{S}(\mathbf{R})$ to the Hamiltonian leaves all down-stream physical observables unchanged. In practice, $\mu$ is chosen as the solution that minimizes the overall loss, as established in Wang et al. (2024c). This removes the gauge freedom inherent in the Hamiltonian representation, facilitating more stable and efficient convergence of the neural network.
@@
$\begin{split}\mathrm{loss}(\mathbf{k})&=\mathbb{E}_{\mathbf{k}}\Big[\lambda_{\mathcal{P}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PP}(\mathbf{k}),\widetilde{\mathbf{H}}_{PP}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{\mathcal{Q}}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{QQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{QQ}^{gt}(\mathbf{k},\mu)\big)\\ &\quad+\lambda_{PQ}\cdot\mathrm{MSE}\big(\widetilde{\mathbf{H}}_{PQ}(\mathbf{k}),\widetilde{\mathbf{H}}_{PQ}^{gt}(\mathbf{k},\mu)\big)\Big],\end{split}$ (25)
where $\lambda_{\mathcal{P}},\lambda_{\mathcal{Q}},\lambda_{PQ}$ are tunable hyper-parameters controlling the relative importance of the three terms.
@@
where the value of $\mu$ is determined by $\frac{\partial\mathrm{loss}_{all}}{\partial\mu}=0$. It can be solved analytically by:
@@ Appendix H Dataset Details
The crystal structures were obtained from the Materials Project database, from which a total of approximately 17,000 nonmagnetic compounds were randomly selected. Among them, 12,000 structures were used for training, 2,000 for validation, and 3,000 for testing. The statistical distributions of atomic species and atomic counts in the training, validation, and test sets are illustrated in Figures 5 and 6. Furthermore, the occurrence frequencies of different elements across the three subsets are presented in Figures 7(a)7(c).
@@ Appendix I Implementation Details
[…]
@@
Appendix J Band Structure Results
@@
Appendix K Efficiency Comparison between Our Method and DFT
@@
Ablation@TraceGrad: In this ablation term, we remove the TraceGrad mechanism. Concretely, the supervision from the trace quantity is omitted in the loss function, and the gradient-based mechanism that delivers non-linearity from O(3)-invariant features $z^{(\text{edge})}_{ab}$ to induce O(3)-equivariant features via $\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{\prime\text{(edge)}}_{ab}}$ is also discarded.
@@
We train all of the ablation terms under the same number of epochs, optimizer, and scheduler as the full method (see Appendix I), then evaluate them on the testing set. The $R$-space errors are summarized in Table 4. Beyond $R$-space, because $k$-space is directly tied to downstream quantities (e.g., band structures), we visualize band predictions for Ablation@Loss-k (R-space only), Ablation@Loss-PQ (setting $\lambda_{PQ}=0$) versus the Full Method in Fig. 15.
From Table 4, the Full Method achieves the lowest Gauge MAE. The Full Method reduces the error by $17.6\%$, $52.3\%$, $20.7\%$, $23.8\%$, $12.2\%$, and $5.2\%$ compared with Ablation@Input, Ablation@Output, Ablation@TraceGrad, Ablation@Ensemble, Ablation@Loss-k, and Ablation@Loss-PQ, respectively. As shown in Fig. 15, Ablation@Loss-k, which removes the $k$-space supervision and relies solely on real-space loss, produces band structures with frequent ghost states: in many cases, while most $k$-points are predicted reasonably well, some $k$-points exhibit abrupt and severe deviations from the ground truth—hallmarks of non-physical artifacts. This phenomenon mainly arises from the error amplification effect analyzed in Appendix F, where the large condition number of the overlap matrix can magnify small real-space errors into significant $k$-space deviations. Importantly, such sparse but catastrophic failures cannot be effectively captured by real-space loss alone. Ablation@Loss-PQ, which augments the training with $k$-space supervision on the intra-subspace blocks ($P$ and $Q$), demonstrates better performance than Ablation@Loss-k, but still fails to completely suppress ghost states. The reason is that unphysical couplings between the low-energy subspace $P$ and the high-energy subspace $Q$ remain unpenalized, and these couplings are precisely what give rise to unphysical artifacts in the band structures. In contrast, the Full Method introduces an important penalty on the $PQ$ cross block, which has clear physical significance: for the exact Hamiltonian, $P$ and $Q$ are strictly decoupled, and any spurious $PQ$ couplings in the predicted Hamiltonian are the direct source of ghost states. By explicitly enforcing this decoupling, the $PQ$ loss term addresses the root cause of the artifacts. As a result, the full method produces band structures in excellent agreement with first-principles DFT and free of ghost states. This comparison clearly demonstrates the necessity of our $k$-space loss design, in particular the $PQ$ penalty, for ensuring the physical reliability of predicted band structures.
@@
Appendix M Comparison with Related Work
In the field of electronic-structure Hamiltonian prediction for periodic material systems, one of the most representative methods is DeepH-E3 (Gong et al., 2023). As a pioneering effort, DeepH-E3 has played a crucial role in introducing equivariant networks to the calculation of electronic structures of materials. We aim to perform a clear comparison with it to demonstrate the innovations and advantages of our approach compared with DeepH-E3.
First, DeepH-E3 uses randomly initialized element (node) and element pair (edge) embeddings trained from scratch as input descriptors. However, the sparse nature of these embeddings poses a problem. Since elements are not uniformly distributed in nature, many element pairs appear infrequently, resulting in poorly trained embeddings that struggle with generalization. What’s more, this leads to the need to maintain an embedding cost of $\mathcal{O}(M)$ and $\mathcal{O}(M^{2})$ for nodes and edges, respectively, where $M$ is the number of elements included in the dataset. As the number of elements increases, this cost grows rapidly, leading to significant memory overhead.
In contrast, our approach addresses these issues by replacing the node and edge embeddings with the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, which eliminates the space complexity and sparsity problems. $\mathbf{H}^{(0)}$ captures crucial information about the system’s electronic structure, embedding the fundamental characteristics of different elements into a unified representation space, thereby offering a richer physical context. This enables robust generalization across a wide range of material systems, making our approach better suited for constructing a truly universal model. What’s more, the space complexity of our input descriptor is independent of the number of distinct elements in the dataset, thus avoiding the exponential cost of embeddings. This use of an easily accessible DFT initial state tensor as a descriptor highlights an emerging direction in AI for Science: embedding physically grounded information directly into machine learning models, which enhances predictive power and provides a general principle that can be broadly applied across scientific ML tasks beyond electronic-structure prediction.
Second, the architecture of DeepH-E3 is relatively simple and lacks sufficient expressive power for building a universal model, which is reflected in two aspects. First, it uses a non-attentive structure, with feature fusion performed through simple average pooling. This may not be well-suited to handle highly complex, diverse structures. Second, it only employs the gating mechanism as the non-linear activation function, without fully exploring the non-linear expressiveness. In contrast, our approach builds upon the TraceGrad paradigm with advanced framework to effectively induce non-linearity from invariant quantities and representations, and develops an E(3)-symmetry Transformer architecture with high non-linear expressiveness. This architecture is capable of dynamically weighting features to adapt to different elements and geometric configurations, making it more suitable for building universal large-scale models.
In addition, our network adopts a delta-learning approach, predicting $\Delta H=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$ instead of the Hamiltonian itself. This significantly reduces the complexity of the regression target, making it easier for the network to finely fit and generalize complex systems, especially those with heavy atoms and a large number of orbitals. In contrast, DeepH-E3 directly predicts the entire Hamiltonian, which becomes challenging when dealing with more complex scenarios.
Third, DeepH-E3 only uses the R-space loss function for training, while we propose a joint loss function combining both R-space and k-space. As shown in Section 4.3 and Appendix L, the inclusion of k-space loss improves the physical reliability of downstream quantities, such as band structure, and prevents ghost states, which appear as abrupt discontinuities at isolated $\mathbf{k}$-points, caused by the amplification of small R-space errors in k-space (Appendix F). By addressing these issues, our method is better suited for handling a wide range of materials and elements, making it more adaptable and reliable across various scientific applications.
We apply DeepH-E3 to our Materials-HAM-SOC dataset, and follow the same training-validation-testing pipeline on our dataset as described in Appendix H. In the feature layers near the output layer and the output itself, we construct tensor representations that correspond to our dataset’s atomic orbital basis sets, which extend to 4s2p2d1f, while leaving the rest of the DeepH-E3 unchanged. These modifications make the comparison more relevant and fair. Unfortunately, even with these adjustments, DeepH-E3 achieves a high R-space error on our testing set, failing to converge to a reasonable error range, as shown in Table 5. In contrast, our method achieves a low error value. This result confirms that DeepH-E3 is not suited for the task of universal electronic structure Hamiltonian prediction, as we have previously analyzed. The research in the DeepH-E3 paper focuses on specialized scenarios, where the training and testing sets consist of perturbations from molecular dynamics simulations of the same material. In contrast, our task involves predicting the Hamiltonian across a broad range of materials, which requires a more generalizable approach that DeepH-E3’s specialized design cannot handle effectively.
Furthermore, to compare the performance in specialized scenarios, we conduct experiments on the Monolayer Graphene (MG) and Monolayer MoS2 (MM) datasets, which are released by the DeepH series (Li et al., 2022; Gong et al., 2023), as introduced in Table 6.
Table 5: Comparison of R-space errors for DeepH-E3 and our method on the testing set of Materials-HAM-SOC. All values are in meV.
Method Gauge MAE
DeepH-E3 12.605
Our Method 1.417
Table 6: Overview of the Monolayer Graphene (MG) and Monolayer MoS2 (MM) datasets. $m$: number of samples in the current dataset; $a$: number of atoms per unit cell in the current dataset.
Statistic Types MG $M$
Training $m$ 270 300
$a$ 72 75
Validation $m$ 90 100
$a$ 72 75
Testing $m$ 90 100
$a$ 72 75
For both datasets, we use the same training, validation, and testing sets as those used in DeepH-E3. To ensure fairness, we retrain our method under identical conditions, without pre-training, and measure errors using the classic MAE (Mean Absolute Error) metric, in alignment with DeepH-E3. In addition to DeepH-E3, we also compare our method with the original TraceGrad work (Yin et al., 2025), which extends the DeepH-E3 backbone network by constructing non-linear equivariant representations from invariant ones.
[…]
Dataset MAE
[…]
MG 0.251 0.175 0.102
$M$ 0.406 0.285 0.163
These results show that our method, even after cutting out some modules, still dramatically outperforms both DeepH-E3 and the TraceGrad work in prediction accuracy for these material systems. While the original TraceGrad work explores effective methods for constructing non-linear equivariant features, it remains a simple graph network based on average pooling. In contrast, our approach extends the TraceGrad paradigm as a more advanced E(3)-symmetry Transformer architecture that can dynamically weight and adapt non-linear equivariant features, which is more expressive than the simple graph network. In addition to demonstrating strong generalization across a variety of materials in previous sections, these results show that our method excels in specialized systems as well, making it a powerful tool for both broad and focused applications in computational materials science.
Statements Regarding the Usage of LLMs
No large language models were used for this work.